OpenStudy (anonymous):

I WILL MEDAL The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below: f(x) = 0.69(1.03)x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer. (5 points) Part B: The table below shows the price f(t), in dollars, of product B after t years: t (number of years) 1 2 3 4 f(t) (price in dollars) 10,100 10,201 10,303.01 10,406.04

OpenStudy (anonymous):

Which product recorded a greater percentage change in price over the previous year?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

@Kainui @ganeshie8

OpenStudy (anonymous):

please help me @sammixboo @Hero

OpenStudy (anonymous):

@heretohelpalways please help

OpenStudy (anonymous):

I only need part B and c (help with), because a is: 12500(0.82)=10250, so it is decreasing by 2250, right?

OpenStudy (anonymous):

@undeadknight26 @TheSmartOne

OpenStudy (anonymous):

please helllppp Wait, I need a too, because i dont know what percent?

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

@ali2x2

OpenStudy (anonymous):

what is 0.69(1.03)

OpenStudy (anonymous):

.7107

OpenStudy (anonymous):

what happens if you multiply any number by something smaller than 1?

OpenStudy (ali2x2):

k

OpenStudy (ali2x2):

PART I You know that this function is DECREASING because 0.63, the number inside of the parenthesis, is LESS THAN THE NUMBER 1. I [think I] know the way to determine by what percentage it is decreasing. [I think] My teacher taught me that I have to subtract the number, in this case 0.63, from 1. So, 1 - 0.63 = 0.37. So it is decreasing by 0.37.

OpenStudy (anonymous):

this thing is pretty intuitive

OpenStudy (ali2x2):



OpenStudy (ali2x2):

part B

OpenStudy (ali2x2):

im sorta confused sorry ;c ill do a google search

OpenStudy (ali2x2):

lmao or mr.doge here can help?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what just happened..

OpenStudy (anonymous):

you got the first part right yet?

OpenStudy (anonymous):

ehhhh I am not sure, as of right now i have that it is decreasing by 37%

OpenStudy (anonymous):

0.69(1.03)=7.107 1-7.107=28.93%

OpenStudy (anonymous):

.69*1.03 is .7107

OpenStudy (anonymous):

you dont really need to care about whats inside or outside, multiplication is commutative

OpenStudy (anonymous):

yea f(x) = 0.69 * 1.03 * x

OpenStudy (anonymous):

you get it yet?

OpenStudy (anonymous):

The x is an exponent, so it is 1.03*.69^x..

OpenStudy (anonymous):

next time put that ^ sign

OpenStudy (anonymous):

lol sorry I didnt pick up that it didnt show till now, sorry.

OpenStudy (anonymous):

the function is actually increasing in that case

OpenStudy (anonymous):

its an exponential curve

OpenStudy (anonymous):

THE VALUE IN THE PARENTHESIS IS MORE THAN ONE. SO YAHHHH ITS INCREASING.

OpenStudy (anonymous):

so confused omfg

OpenStudy (anonymous):

regard the %, try plugging in x=0 and x=1 into the function

OpenStudy (anonymous):

f(x)=.69(1.03)^0 .69*1 .69, or 69%

OpenStudy (anonymous):

f(x)=.69(1.03)^1 .69* 1.03 .7107

OpenStudy (anonymous):

ok now take\[f(1)-f(0)\]

OpenStudy (anonymous):

.7107-.69

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

.0207

OpenStudy (anonymous):

what % is 0.0207 of 0.69?

OpenStudy (anonymous):

.00014283

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

I am not sharp. Can you help?

OpenStudy (anonymous):

\[\frac{ 0.0207 }{ 0.69 } \times 100\]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3% increase

OpenStudy (anonymous):

part b is easy

OpenStudy (anonymous):

take year 2 - year 1

OpenStudy (anonymous):

okayy. so 101

OpenStudy (anonymous):

now find the percentage like what you just did for the previous part

OpenStudy (anonymous):

101%?

OpenStudy (anonymous):

no 10.1%?

OpenStudy (anonymous):

\[\frac{ 10201-10100 }{ 10100 } \times 100\]

OpenStudy (anonymous):

1%

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now the answer for part b is in front of you compare % of product a & b

OpenStudy (anonymous):

okay so product a: 3%, product b=1%?

OpenStudy (anonymous):

yes, and which is greater?

OpenStudy (anonymous):

PRODUCT A?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

sorry about caps

OpenStudy (anonymous):

np